OpenStudy (smileyxl3):

For the sequence 15, 22, 29, 36, 43, ... , solve the following problems

OpenStudy (smileyxl3):

a. Write a recursive formula for the sequence. b. Write an explicit formula for the sequence. c. Find the sum of the first 100 terms of the sequence.

OpenStudy (anonymous):

For a. I would say \[ a_1=15 \quad \text{and}\quad a_{n+1} = a_n +7 \]

OpenStudy (rane):

whts the rule or difference have u found out for the numbers above ? @smileyxl3

OpenStudy (nirmalnema):

a= first term of the series= 15 d = common difference=a2=a1 = 7

OpenStudy (nirmalnema):

sum of 1st hundred terms =100/2 (2(15)+(100-1)7) 50(30+(99)7) = 36150

OpenStudy (smileyxl3):

@nirmalnema I get everything except why you have to multiply 2*15

hartnn (hartnn):

because the formula says so, \(\Large S_n= (n/2) (\color {red}2a_1+(n-1)d)\)

OpenStudy (smileyxl3):

@hartnn I have no idea why I didn't see that in the first place. Thanks

hartnn (hartnn):

no problem, welcome ^_^